OpenStudy (anonymous):

Please answer both questions 1) a fair coin and a six sided die are tossed stimutaneously. What is the probability of obtaining jointly: a.a head and a 3 B.a head and an even number C.a tail and a number less than 5 D.a tail and a number greater than 4 2) two fair coins are tossed. What is the probability that both land heads up?

OpenStudy (anonymous):

1) 1/2*1/6=A 1/2*1/2=B 1/2*2/3=C 1/2*1/3=D 2) 1/2*1/2=#2

OpenStudy (campbell_st):

P(head) = 12/ P(3) = 1/6 P(head & 3) = 1/2 * 1/6

OpenStudy (campbell_st):

basically find the probability of each event then multiply to get the answer.